Adult tickets for the play cost $5 each and student tickets cost $3 each. A total of 720 tickets worth of $3,120 were sold. How many student tickets were sold?
Answer by mananth(16946)

Adult tickets $5 each ----------- x tickets
student tickets cost $3 each------- y tickets
total of 720 tickets
worth of $3,120 were sold.
x+y = 720 .............1
5x + 3y= 3120 .............2

Eliminate y
multiply (1)by -3
Multiply (2) by 1
-3 x -3 y = -2160
5 x 3 y = 3120
Add the two equations
2 x = 960
/ 2
x = 480
plug value of x in (1)
1 x + 1 y = 720
480 + 1 y = 720
1 y = 720 -480
1 y = 240
y = 240
adults 480 tickets
students 240 tickets
m.ananth@hotmail.ca